The formula for calculating the sum of all natural integers from 1 to n is well-known:


Sn = 1 + 2 + 3 + ... + n =


n


2 + n


2


Similary, we know about the formula for calculating the sum of the first n squares:


Qn = 1 · 1 + 2 · 2 + 3 · 3 + ... + n · n =


n


3


3


+


n


2


2


+


n


6


Now, we reduce one of the two multipliers of each product by one to get the following sum:


Mn = 0 · 1 + 1 · 2 + 2 · 3 + 3 · 4 + ... + (n − 1) · n


Find an explicit formula for calculating the sum Mn.

1
Expert's answer
2022-02-21T18:31:02-0500

Here "M_n=0.1+1.2+2.3+3.1+....+(n-1)n"


"n^{th}" term is-

"t_n=(n-1)n\n\n =n^2-n"


So The Sum ,


"M_n=\sum t_n=\sum [n^2-n]=\sum n^2-\sum n"


"=\dfrac{n}{6}(n+1)(2n+1)-\dfrac{n}{2}(n+1)\\=\dfrac{n}{2}(n+1)[\dfrac{2n+1}{3}-1]\\=\dfrac{n}{6}(n+1)(2n-2)\\=\dfrac{n}{3}(n^2-1)"